ok its a max height question and im unsure of how to go about solving it.
the path of a basketball shot can be modelled by the equation h= -0.125d^2 +2.5 where h is the height of the ball and d is the horizontal distance.
find the max height
find the horizontal distance of the ball when it reaches max height
and how far from the floor is the ball when the player releases it ?
thats all thanks

Answer by stanbon(75887)

The path of a basketball shot can be modeled by the equation
h(d) = -0.125d^2 +2.5 where h is the height of the ball and d is the horizontal distance.
Find the max height.
Find the horizontal distance of the ball when it reaches max height
and how far from the floor the ball is when the player releases it ?
---------------------------------------------------------------------------
h(d) = -0.125d^2 + 2.5
----
max height occurs when d = -b/2a = -2.5/(2*-0.125) = 10 ft.
--------
height of the ball when the player releases it:
h(0) = -0.125(0)^2+2.5 = 0 + 2.5 = 2.5
==========================================
Cheers,
Stan H.